OpenStudy (fellowroot):

In xyz-space, an equation of the plane tangent to the surface z=sin(y)e^(-x) at the point where x=0 and y=(π/2) is (A) x+y=1 (B) x+z=1 (C) x−z=1 (D) y+z=1 (E) y−z=1

OpenStudy (anonymous):

did you find the gradient?

OpenStudy (anonymous):

the gradient is normal to the surface, and so, normal to the tangent plane...

OpenStudy (anonymous):

you there?

OpenStudy (fellowroot):

I'll try the gradient and see what I get.

OpenStudy (anonymous):

\[<-\sin y*e^{-x} i + \cos y *e^{-x} j - k>\]

OpenStudy (anonymous):

for x=1, y=pi/2 it is <-i -k>